# Refined Risk Bounds for Unbounded Losses via Transductive Priors

### v1

@@ 5.2.2 Aggregating the slab-experts
To further improve the implementability of Algorithm 1 with generalized slab-experts $\mathcal{A}^{+}$ instead of the slab-experts $\mathcal{A}$, we can replace the enumeration of the generalized slab-experts with sampling. Concretely, we first sample a random subset of at most $d+1$ points of the covariate set ${K}$. Next, we sample the labels randomly. Finally, we apply SVM to obtain a halfspace. Since any halfspace is sampled with probability at least $\Omega(\mathrm{poly}(1/T^{d}))$, the optimal slab-expert $B_{{\theta^{\star}},\alpha}$ will also be sampled with probability at least $\Omega(\mathrm{poly}(1/T^{d}))$. Thus, the regret can be upper bounded by $O(d\log T)$. The advantage of such a sampling scheme is that, for a set of covariates in general position (no $d$ points are linearly dependent), random labeling of any $d+1$ subset will typically lead to a halfspace. This is validated by Cover (1965, Theorem 1). Specifically, for any $d+1$ points in general position, the number of distinct halfspaces is at least $2\sum\limits_{k=0}^{d-1}\binom{d}{k}=2^{d+1}-2\geq 2^{d}$ according to Cover (1965, Theorem 1). However, there are at most $2^{d+1}$ different labeling available. Thus, random labeling allows for a halfspace separation with probability at least $1/2$. Moreover, since the aggregation follows the exponentially weighted averaging
$\tilde{\rho}_{t}(x_{t},\cdot)=\sum\limits_{B\in\mathcal{A}^{+}}p_{t}(B)B(x_{t},\cdot),\text{~{}~{}where~{}~{}}p_{t}(B)\propto\exp\left(\sum\limits_{i=1}^{t-1}\log B(x_{i},y_{i})\right).$
The distribution $p_{t}$ on the set of generalized slab-experts $\mathcal{A}^{+}$ is a Gibbs-type distribution that can be represented as the stationary distribution of a Markov chain generated by the Metropolis-Hastings (MH) algorithm (Robert and Casella, 1999, Section 7.3), in a manner similar to that in (Rigollet and Tsybakov, 2012, Section 7).
@@ 6 Online-to-Batch Conversion: Fixed Design vs Random Design
In the problem of statistical learning, the learner is given a sample $S_{T}=\{(X_{1},Y_{1}),\ldots,(X_{T},Y_{T})\}$ consisting of i.i.d. pairs sampled according to some unknown distribution $\mathcal{D}$ on the space $\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ is the covariate space, and $\mathcal{Y}$ is the observation space. The learner is further given a function class $\mathcal{F}:\mathcal{X}\to\mathcal{Y}$ and a loss function $\ell(\cdot,\cdot):\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$. The learner’s goal is to output a (possibly randomized) predictor $\widehat{f}_{T}$ to minimize the (expected) excess risk defined as
@@
Proposition 6.1 (Transductive online-to-batch conversion).
@@
Now, given a sample $S_{T}$ of size $T$ and an additional test point $(X,Y)$, both drawn i.i.d. from an unknown distribution $\mathcal{D}$, construct the sequence $\widehat{f}_{1},\ldots,\widehat{f}_{T+1}$ , where $\widehat{f}_{t+1}$ is based on the labeled sample $(X_{1},Y_{1}),\dots,(X_{t},Y_{t})$ and the unlabeled set $\{X_{1},\ldots,X_{T},X\}$ [^3] Accordingly, we assume that at the level of the entire unlabeled dataset, the order of the observations does not influence the output of the algorithm, a condition consistently met in our applications. and satisfies the deterministic regret bound (41) with an upper bound on the regret $\operatorname{Reg}_{T+1}^{\operatorname{trd}}$. The final predictor, defined as $\widetilde{f}=\frac{1}{T+1}\sum_{t=1}^{T+1}\widehat{f}_{t}$, satisfies the following excess risk bound:
@@
Proof of Proposition 6.1 Denote $(X_{T+1},Y_{T+1})=(X,Y)$. To highlight the dependence of any predictor $f_{t}$ on both the unlabeled set $\{X_{s}\}_{s=1}^{T+1}$ and the labeled sequence of observations $X_{1:t-1},Y_{1:t-1}$, we use the following notation for the loss $\ell(\widehat{f}_{t}(X_{t}\mid X_{1:t-1},Y_{1:t-1},\{X_{s}\}_{s=1}^{T+1}),Y_{t})$ at the observation $X_{t}$, where $X_{1:t-1}$ denotes the sequence $X_{1},\ldots,X_{t-1}$ and $Y_{1:t-1}$ stands for the sequence $Y_{1},\ldots,Y_{t-1}$, respectively. We first break the loss into summands of conditional expectations with Jensen’s inequality:
@@
=$\stackrel{{\scriptstyle(*)}}{{=}}\mathbb{E}_{X_{1},...,X_{T},X_{T+1}}\left[\mathbb{E}_{(Y_{s}|X_{s})_{s=1}^{t}}\left[\ell(\widehat{f}_{t}(X_{t}|X_{1:t-1},Y_{1:t-1},\{X_{s}\}_{s=1}^{T+1}),Y_{t})\right]\right]$
@@
where for ()(*)( ∗ ), we use the fact that the distribution is not changed with the exchange. Then,
@@
$\leq\mathbb{E}_{S_{T}}\mathbb{E}_{(X_{T+1},Y_{T+1})}\left[\sum\limits_{t=1}^{T+1}\ell(\widehat{f}_{t}(X_{t}),Y_{t})]-\inf_{f\in\mathcal{F}}\sum\limits_{t=1}^{T+1}\ell(f(X_{t}),Y_{t})\right]$
@@ 6.1 Applications of online-to-batch conversion
In this section, we demonstrate two applications of the online-to-batch conversion result presented in Proposition 6.1.
@@
Consider the regression problem with squared loss. Let the covariates $X_{1},\dots,X_{T}$ and response variables $Y_{1},\dots,Y_{T}$ be i.i.d. copies of the random pair $(X,Y)$ drawn from an unknown distribution. Denote the dataset as $S_{T}=\{(X_{1},Y_{1}),\dots,(X_{T},Y_{T})\}$. We make no assumptions on the distribution of $X\in\mathbb{R}^{d}$, but assume that $|Y|\leq m$ almost surely. By combining the regret bound from Theorem 2.1 (with $\lambda=1/T$) and the online-to-batch conversion argument from Proposition 6.1, we immediately construct a computationally efficient batch predictor $\widetilde{f}$, based on $S_{T}$, that satisfies the following excess risk bound:
@@
$\mathbb{E}_{S_{T}}\left[-\log(\widetilde{p}(X,Y))-\inf_{\theta\in\mathbb{R}^{d}}\mathbb{E}\left(-\log(\sigma(Y\langle X,\theta\rangle))\right)\right]\leq\frac{11d\log(e(T+1))}{T+1}.$
@@
$\mathbb{E}_{S_{T}}\mathbb{E}_{X}\mathcal{KL}\left(\mathrm{Ber}(\sigma(\langle X,\theta^{\star}\rangle))\|\mathrm{Ber}(\widetilde{p}(X,1))\right)\leq\frac{11d\log(e(T+1))}{T+1}.$
This result differs from previous excess risk bounds for logistic regression found in the literature (see, e.g., (Bach, 2010; Ostrovskii and Bach, 2021; Foster et al., 2018; Mourtada and Gaïffas, 2022; Vijaykumar, 2021; Bilodeau et al., 2023; van der Hoeven et al., 2023)) in that it imposes no assumptions on the marginal distribution of the covariates, and the regret bound does not depend on the norm of the minimizer $\theta^{\star}$.
Proof of Corollary 6.1 The proof follows directly from combining Theorem 3.2 with Proposition 6.1. The second statement follows from the standard computation, showing that in the well-specified case, the logarithmic loss excess risk corresponds to the Kullback-Leibler divergence between the corresponding distributions. ∎
@@ 6.2 Polynomial time algorithm for batch logistic regression
We will present the proof for the distribution $\mathcal{D}_{\theta^{\star}}$ where the covariate $X$ follows a uniform distribution on the scaled sphere $\sqrt{d}\mathcal{S}^{d-1}$. Nevertheless, our proof only requires mild assumptions (Assumptions 6.1 and 6.2) on the marginal distribution of the covariates $X$. In effect, our proof extends to and well beyond the standard Gaussian distribution.
@@
Definition 6.1 (Clipped linear separator).
@@
$k_{\theta}(x,y)=\frac{1}{2T}+\left(1-\frac{1}{T}\right)\operatorname{sign}(y\langle x,\theta\rangle).$
Namely, the expert assigns probability $1-1/(2T)$ to the label predicted by the linear separator and $1/(2T)$ to the alternative one.
@@
      Let the prediction be the aggregation on the slab-experts $\tilde{\rho}_{t}(\cdot,\cdot)=\sum\limits_{B\in\mathcal{A}}p_{t}(B)B(\cdot,\cdot)\text{~{}~{}where~{}~{}}p_{t}(B)\propto\exp\left(\sum\limits_{i=1}^{t-1}-\log B(X_{i},Y_{i})\right)$ (42) and $B(X_{i},Y_{i})$ is probability of $B$ predicting $Y_{i}$ on $X_{i}$.
@@
There are two cases to analyze. The first case assumes that the norm of $\theta^{*}$ is bounded by $T^{4}$. In this scenario, by setting $\alpha<1/T^{9}$, the expert using the transductive prior achieves a rate of $O(d\log T/T)$ (Proposition 3.1).
In the second case, where the norm of $\theta^{*}$ exceeds $T^{4}$, we further divide the analysis into three subcases: 1. The first subcase is when the data is not separable by $\theta^{*}$, which occurs with small probability due to the large norm of $\theta^{*}$ (Lemma 6.1). 2. The second subcase is when $\widehat{\theta}$ deviates from $\theta^{*}$ by more than $(d+\log T)/T$, which also happens with small probability (Lemma 6.2). 3. The final subcase occurs when $\widehat{\theta}$ is close to $\theta^{*}$, in which case the clipped linear separator with $\widehat{\theta}$ achieves a rate of $O((d+\log T)\log T/T)$ (Lemma 6.3). This completes the proof.
We provide a proof for covariates distributed uniformly on the scaled sphere $\sqrt{d}\mathcal{S}^{d-1}$, relying on two key properties of the covariate distribution, as described in Assumptions 6.1 and 6.2. These properties also hold when the covariates follow a standard Gaussian distribution, thus extending our proof to that case as well.
Assumption 6.1 (Bounded eigenvalue).
@@
The choice of the constant one in the upper bound $\lambda_{\max}\left(\mathbb{E}[XX^{\top}]\right)\leq 1$ can be relaxed to any constant and will appear as a log factor in the final risk bound. Both the uniform distribution on the sphere and the standard Gaussian distribution satisfy the bounded eigenvalue assumption (Assumption 6.1). For the former, we have for any $\theta$ on the scaled sphere $\sqrt{d}\mathcal{S}^{d-1}$:
@@
Assumption 6.2 (Small probability in thin slabs).
@@
The choice of the constant $\sqrt{d}/\sqrt{\pi}$ can be relaxed to any polynomial dependence on $d$ and will appear as a log factor in the final risk bound. Both the uniform distribution on the scaled sphere $\sqrt{d}\mathcal{S}^{d-1}$ and the standard Gaussian distribution satisfy Assumption 6.2. For the former, if $d=1$, $\langle X,\theta\rangle=\pm 1$, thus
@@
[…]
where the first equality is due to the formula for the projection of a uniform distribution on the interval $[-1,1]$, and the inequality follows from the fact that the density is upper bounded by $\frac{d}{2\sqrt{\pi}}$. For covariates $X$ distributed according to the standard Gaussian, Assumption 6.2 also holds up to a constant factor since the projection in any direction is standard normal with density upper bounded by $1/\sqrt{2\pi}$.
With Assumption 6.2, when $\|\theta^{\star}\|$ is large, then with high probability, the data generated will be separable as shown by the following Lemma 6.1.
Lemma 6.1.
Let $T>1$ be an integer. For any $\theta^{\star}\in\mathbb{R}^{d}$, with $\|\theta^{\star}\|>T^{4}$, with probability at least $1-d/T$, the set $\{(X_{t},Y_{t})\}_{t=1}^{T}$ drawn i.i.d. from $\mathcal{D}_{\theta^{\star}}$ satisfies $Y_{t}=\operatorname{sign}(\langle X_{t},\theta^{\star}\rangle)$ for all $t$.
Proof of Lemma 6.1 If $d\geq T^{2}$, we have nothing to prove. Thus, we assume $d<T^{2}$ for the following.
@@
$\mathbb{P}\left(\forall t,Y_{t}\langle X_{t},\theta^{\star}\rangle>0\right)=\left(\mathbb{P}\left(Y_{1}\langle X_{1},\theta^{\star}\rangle>0\right)\right)^{T}.$
Furthermore, by the conditional dependence of $Y_{1}$ on $X_{1}$, we have
$\mathbb{P}\left(Y_{1}\langle X_{1},\theta^{\star}\rangle>0\right)$ $\geq\mathbb{P}\left(Y_{1}\langle X_{1},\theta^{\star}\rangle>0,\left|\langle X_{1},\theta^{\star}\rangle\right|>T\right)$
$\geq\mathbb{P}\left(\left|\langle X_{1},\theta^{\star}\rangle\right|>T\right)\cdot\frac{1}{1+e^{-T}}.$
Then, by Assumption 6.2, we have
$\mathbb{P}\left(\left|\langle X_{1},\theta^{\star}\rangle\right|>T\right)$ $=1-\mathbb{P}\left(\left|\langle X_{1},\theta^{\star}\rangle\right|\leq T\right)$
$\geq 1-\mathbb{P}\left(\left|\langle X_{1},\theta^{\star}/\|\theta^{\star}\|\rangle\right|\leq 1/T^{3}\right)$
@@
$\mathbb{P}\left(\forall t,Y_{t}\langle X_{t},\theta^{\star}\rangle>0\right)$ $\geq\left(\left(1-\frac{\sqrt{d}}{\sqrt{\pi}T^{3}}\right)\cdot\frac{1}{1+e^{-T}}\right)^{T}$
@@
Whenever the data $\{(X_{t},Y_{t})\}_{t\in[T]}$ is separable, the hard margin SVM will, with high probability, output a classifier $\hat{\theta}$ where the disagreement region for the linear separators induced by $\hat{\theta}$ and $\theta^{\star}$ is only $O\left(\frac{d+\log T}{T}\right)$, according to the recent result of Bousquet et al. (2020). This is concretized by the following Lemma 6.2.
Lemma 6.2.
Let $T>1$ be an integer. For any ${\theta^{\star}}\in\mathbb{R}^{d}$, with $\|{\theta^{\star}}\|>T^{2}$, suppose the $\{X_{t}\}_{t=1}^{T}$ are drawn i.i.d. from the distribution $\mathcal{D}_{{\theta^{\star}}}$ conditioning on the event $\{Y=\operatorname{sign}(\langle X,{\theta^{\star}}\rangle)\}$. Then, conditioning on the event $\{\forall t,Y_{t}=\operatorname{sign}(\langle X_{t},{\theta^{\star}}\rangle)\}$, with probability at least $1-d/T$, the $\mathrm{SVM}$ solution $\hat{\theta}$ on $\{(X_{t},Y_{t})\}_{t\in[T]}$ achieves
$\hat{\theta}\in\operatorname*{arg\,min}_{\theta}\sum\limits_{t=1}^{T}\mathbbm{1}(Y_{t}\langle X_{t},\theta\rangle<0)$
@@
Proof of Lemma 6.2 If $d>T$, then the argument automatically holds. Thus, we assume hereafter that $d\leq T$. This proof is a revised version of Hsu and Mazumdar (2024, Lemma 14). For any $\theta$, the error of the linear separator defined by $\theta$ in $0/1$ loss is defined as
$\mathrm{err}_{{\theta^{\star}}}(\theta)$ :$\vcentcolon=\mathbb{E}\left[\mathbbm{1}(Y\langle X,\theta\rangle<0)\right]$

### v3

@@ 5.2.2 Aggregating the slab-experts
To further improve the implementability of Algorithm 1 with generalized slab-experts $\mathcal{A}^{+}$ instead of the slab-experts $\mathcal{A}$, we can replace the enumeration of the generalized slab-experts with sampling. Concretely, we first sample a random subset of at most $d+1$ points of the covariate set K{K}. Next, we sample the labels randomly. Finally, we apply SVM to obtain a halfspace. Since any halfspace is sampled with probability at least $\Omega(\mathrm{poly}(1/T^{d}))$, the optimal slab-expert $B_{{\theta^{\star}},\alpha}$ will also be sampled with probability at least $\Omega(\mathrm{poly}(1/T^{d}))$. Thus, the regret can be upper bounded by $O(d\log T)$. The advantage of such a sampling scheme is that, for a set of covariates in general position (no $d$ points are linearly dependent), random labeling of any $d+1$ subset will typically lead to a halfspace. This is validated by Cover (1965, Theorem 1). Specifically, for any $d+1$ points in general position, the number of distinct halfspaces is at least $2\sum\limits_{k=0}^{d-1}\binom{d}{k}=2^{d+1}-2\geq 2^{d}$ according to Cover (1965, Theorem 1). However, there are at most $2^{d+1}$ different labeling available. Thus, random labeling allows for a halfspace separation with probability at least $1/2$. Moreover, since the aggregation follows the exponentially weighted averaging
$\tilde{\rho}_{t}(x_{t},\cdot)=\sum\limits_{B\in\mathcal{A}^{+}}p_{t}(B)B(x_{t},\cdot),\text{\leavevmode\nobreak\ \leavevmode\nobreak\ where\leavevmode\nobreak\ \leavevmode\nobreak\ }p_{t}(B)\propto\exp\left(\sum\limits_{i=1}^{t-1}\log B(x_{i},y_{i})\right).$
The distribution $p_{t}$ on the set of generalized slab-experts $\mathcal{A}^{+}$ is a Gibbs-type distribution that can be represented as the stationary distribution of a Markov chain generated by the Metropolis-Hastings (MH) algorithm (Robert and Casella, 1999, Section 7.3), in a manner similar to that in (Rigollet and Tsybakov, 2012, Section 7). Note that this approach provides an implementable sampling procedure rather than offering any computational advantages.
@@ 6 Online-to-Batch Conversion: Fixed Design vs Random Design
In the problem of statistical learning, the learner is given a sample $S_{T}=\{(X_{1},Y_{1}),\ldots,\\ (X_{T},Y_{T})\}$ consisting of i.i.d. pairs sampled according to some unknown distribution $\mathcal{D}$ on the space $\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ is the covariate space, and $\mathcal{Y}$ is the observation space. The learner is further given a function class $\mathcal{F}:\mathcal{X}\to\mathcal{Y}$ and a loss function $\ell(\cdot,\cdot):\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$. The learner’s goal is to output a (possibly randomized) predictor $\widehat{f}_{T}$ to minimize the (expected) excess risk defined as
@@
Proposition 25 (Transductive online-to-batch conversion).
@@
Now, given a sample $S_{T}$ of size $T$ and an additional test point $(X,Y)$, both drawn i.i.d. from an unknown distribution $\mathcal{D}$, construct the sequence $\widehat{f}_{1},\ldots,\widehat{f}_{T+1}$ , where $\widehat{f}_{t+1}$ is based on the labeled sample $(X_{1},Y_{1}),\dots,(X_{t},Y_{t})$ and the unlabeled set $\{X_{1},\ldots,X_{T},X\}$ [^4] Accordingly, we assume that at the level of the entire unlabeled dataset, the order of the observations does not influence the output of the algorithm, a condition consistently met in our applications. Whether this assumption is indeed required for transductive online-to-batch conversions in general remains open. and satisfies the deterministic regret bound (39) with an upper bound on the regret $\operatorname{Reg}_{T+1}^{\operatorname{trd}}$. The final predictor, defined as $\widetilde{f}=\frac{1}{T+1}\sum_{t=1}^{T+1}\widehat{f}_{t}$, satisfies the following excess risk bound:
@@
Proof of Proposition 25 Denote $(X_{T+1},Y_{T+1})=(X,Y)$. To highlight the dependence of any predictor $f_{t}$ on both the unlabeled set $\{X_{s}\}_{s=1}^{T+1}$ and the labeled sequence of observations $X_{1:t-1},Y_{1:t-1}$, we use the following notation for the loss $\ell(\widehat{f}_{t}(X_{t}\mid X_{1:t-1},Y_{1:t-1},\{X_{s}\}_{s=1}^{T+1}),Y_{t})$ at the observation $X_{t}$, where $X_{1:t-1}$ denotes the sequence $X_{1},\ldots\\ ,X_{t-1}$ and $Y_{1:t-1}$ stands for the sequence $Y_{1},\ldots,Y_{t-1}$, respectively. We first break the loss into summands of conditional expectations with Jensen’s inequality:
@@
$\stackrel{{\scriptstyle(*)}}{{=}}\mathbb{E}_{X_{1},...,X_{T},X_{T+1}}\left[\mathbb{E}_{(Y_{s}|X_{s})_{s=1}^{t}}\left[\ell(\widehat{f}_{t}(X_{t}|X_{1:t-1},Y_{1:t-1},\{X_{s}\}_{s=1}^{T+1}),Y_{t})\right]\right]$
@@
where for ()(*), we use the fact that the distribution is not changed with the exchange. Then,
@@
$\leq\mathbb{E}_{S_{T}}\mathbb{E}_{(X_{T+1},Y_{T+1})}\left[\sum\limits_{t=1}^{T+1}\ell(\widehat{f}_{t}(X_{t}),Y_{t})-\inf_{f\in\mathcal{F}}\sum\limits_{t=1}^{T+1}\ell(f(X_{t}),Y_{t})\right]$
@@ 6.1 Applications of online-to-batch conversion
In this section, we demonstrate two applications of the online-to-batch conversion result presented in Proposition 25.
@@
Consider the regression problem with squared loss. Let the covariates $X_{1},\dots,X_{T}$ and response variables $Y_{1},\dots,Y_{T}$ be i.i.d. copies of the random pair $(X,Y)$ drawn from an unknown distribution. Denote the dataset as $S_{T}=\{(X_{1},Y_{1}),\dots,(X_{T},Y_{T})\}$. We make no assumptions on the distribution of $X\in\mathbb{R}^{d}$, but assume that $|Y|\leq m$ almost surely. By combining the regret bound from Theorem 1 (with $\lambda=1/T$) and the online-to-batch conversion argument from Proposition 25, we immediately construct a computationally efficient batch predictor $\widetilde{f}$, based on $S_{T}$, that satisfies the following excess risk bound:
@@
$\mathbb{E}_{S_{T}}\left[-\log(\widetilde{p}(X,Y))-\inf_{\theta\in\mathbb{R}^{d}}\mathbb{E}\left(-\log(\sigma(Y\langle X,\theta\rangle))\right)\right]\leq\frac{6d\log(e(T+1))}{T+1}.$
@@
$\mathbb{E}_{S_{T}}\mathbb{E}_{X}\mathcal{KL}\left(\mathrm{Ber}(\sigma(\langle X,\theta^{\star}\rangle))\|\mathrm{Ber}(\widetilde{p}(X,1))\right)\leq\frac{6d\log(e(T+1))}{T+1}.$
This result differs from previous excess risk bounds for logistic regression found in the literature (see, e.g., (Bach, 2010; Ostrovskii and Bach, 2021; Foster et al., 2018; Mourtada and Ga¨ıffas, 2022; Vijaykumar, 2021; Bilodeau et al., 2023; van der Hoeven et al., 2023)) in that it imposes no assumptions on the marginal distribution of the covariates, and the regret bound does not depend on the norm of the minimizer $\theta^{\star}$.
Proof of Corollary 26 The proof follows directly from combining Theorem 13 with Proposition 25. The second statement follows from the standard computation, showing that in the well-specified case, the logarithmic loss excess risk corresponds to the Kullback-Leibler divergence between the corresponding distributions. ∎
@@ 6.2 Polynomial time algorithm for batch logistic regression
We will present the proof for the distribution $\mathcal{D}_{\theta^{\star}}$ where the covariate $X$ follows a uniform distribution on the scaled sphere $\sqrt{d}\mathcal{S}^{d-1}$. Nevertheless, our proof only requires mild assumptions (29 and 30) on the marginal distribution of the covariates $X$. In effect, our proof extends to and well beyond the standard Gaussian distribution.
@@
Definition 27 (Clipped linear separator).
@@
$k_{\theta}(x,y)=\begin{cases}1-\frac{1}{2T},&y\langle x,\theta\rangle\geq 0,\\ \frac{1}{2T},&\textit{otherwise.}\end{cases}$
Namely, the expert assigns probability 11/(2T)1-1/(2T) to the label predicted by the linear separator and $1/(2T)$ to the alternative one.
@@
   Let the prediction be the aggregation on the slab-experts $\tilde{\rho}_{t}(\cdot,\cdot)=\sum\limits_{B\in\mathcal{A}}p_{t}(B)B(\cdot,\cdot)\text{\leavevmode\nobreak\ \leavevmode\nobreak\ where\leavevmode\nobreak\ \leavevmode\nobreak\ }p_{t}(B)\propto\exp\left(\sum\limits_{i=1}^{t-1}\log B(X_{i},Y_{i})\right)$ (40) and $B(X_{i},Y_{i})$ is probability of $B$ predicting $Y_{i}$ on $X_{i}$.
end for
@@
There are two cases to analyze. The first case assumes that the norm of $\theta^{*}$ is bounded by $T^{4}$. In this scenario, by setting $\alpha<1/T^{9}$, the expert using the transductive prior achieves a rate of $O(d\log T/T)$ (Proposition 15).
In the second case, where the norm of $\theta^{*}$ exceeds $T^{4}$, we further divide the analysis into three subcases: 1. The first subcase is when the data is not separable by $\theta^{*}$, which occurs with small probability due to the large norm of $\theta^{*}$ (Lemma 31). 2. The second subcase is when $\widehat{\theta}$ deviates from $\theta^{*}$ by more than $(d+\log T)/T$, which also happens with small probability (Lemma 32). 3. The final subcase occurs when $\widehat{\theta}$ is close to $\theta^{*}$, in which case the clipped linear separator with $\widehat{\theta}$ achieves a rate of $O((d+\log T)\log T/T)$ (Lemma 33). This completes the proof.
We provide a proof for covariates distributed uniformly on the scaled sphere $\sqrt{d}\mathcal{S}^{d-1}$, relying on two key properties of the covariate distribution, as described in 29 and 30. These properties also hold when the covariates follow a standard Gaussian distribution, thus extending our proof to that case as well.
Assumption 29 (Bounded eigenvalue).
@@
The choice of the constant one in the upper bound $\lambda_{\max}\left(\mathbb{E}[XX^{\top}]\right)\leq 1$ can be relaxed to any constant and will appear as a log factor in the final risk bound. Both the uniform distribution on the sphere and the standard Gaussian distribution satisfy the bounded eigenvalue assumption (29). For the former, we have for any $\theta$ on the scaled sphere $\sqrt{d}\mathcal{S}^{d-1}$:
@@
Assumption 30 (Small probability in thin slabs).
@@
The choice of the constant $\sqrt{d}/\sqrt{\pi}$ can be relaxed to any polynomial dependence on $d$ and will appear as a log factor in the final risk bound. Both the uniform distribution on the scaled sphere $\sqrt{d}\mathcal{S}^{d-1}$ and the standard Gaussian distribution satisfy 30. For the former, if $d=1$, $\langle X,\theta\rangle=\pm 1$, thus
@@
$\mathbb{P}_{X}\left(\left|\langle X,\theta\rangle\right|\leq\varepsilon\right)=\int_{-\varepsilon/\sqrt{d}}^{\varepsilon/\sqrt{d}}\frac{\Gamma(d/2)}{\sqrt{\pi}\Gamma((d-1)/2)}(1-x^{2})^{\frac{d-3}{2}}\,dx\leq\frac{\sqrt{d}\varepsilon}{\sqrt{\pi}},$
where the first equality is due to the formula for the projection of a uniform distribution on the interval [1,1][-1,1], and the inequality follows from the fact that the density is upper bounded by $\frac{d}{2\sqrt{\pi}}$. For covariates $X$ distributed according to the standard Gaussian, 30 also holds up to a constant factor since the projection in any direction is standard normal with density upper bounded by $1/\sqrt{2\pi}$.
With 30, when $\|\theta^{\star}\|$ is large, then with high probability, the data generated will be separable as shown by the following Lemma 31.
Lemma 31.
Let $T>1$ be an integer. For any $\theta^{\star}\in\mathbb{R}^{d}$, with $\|\theta^{\star}\|>T^{4}$, with probability at least 1d/T1-d/T, the set $\{(X_{t}^{\prime},Y_{t}^{\prime})\}_{t=1}^{T}$ drawn i.i.d. from $\mathcal{D}_{\theta^{\star}}$ satisfies $Y_{t}^{\prime}=\operatorname{sign}(\langle X_{t}^{\prime},\theta^{\star}\rangle)$ for all $t$.
Proof of Lemma 31 If $d\geq T^{2}$, we have nothing to prove. Thus, we assume $d<T^{2}$ for the following.
@@
$\mathbb{P}\left(\forall t,Y_{t}^{\prime}\langle X_{t}^{\prime},\theta^{\star}\rangle>0\right)=\left(\mathbb{P}\left(Y_{1}^{\prime}\langle X_{1}^{\prime},\theta^{\star}\rangle>0\right)\right)^{T}.$
Furthermore, by the conditional dependence of $Y_{1}^{\prime}$ on $X_{1}^{\prime}$, we have
$\mathbb{P}\left(Y_{1}^{\prime}\langle X_{1}^{\prime},\theta^{\star}\rangle>0\right)$ $\geq\mathbb{P}\left(Y_{1}^{\prime}\langle X_{1}^{\prime},\theta^{\star}\rangle>0,\left|\langle X_{1}^{\prime},\theta^{\star}\rangle\right|>T\right)$
$\geq\mathbb{P}\left(\left|\langle X_{1}^{\prime},\theta^{\star}\rangle\right|>T\right)\cdot\frac{1}{1+e^{-T}}.$
Then, by 30, we have
$\mathbb{P}\left(\left|\langle X_{1}^{\prime},\theta^{\star}\rangle\right|>T\right)$ $=1-\mathbb{P}\left(\left|\langle X_{1}^{\prime},\theta^{\star}\rangle\right|\leq T\right)$
$\geq 1-\mathbb{P}\left(\left|\langle X_{1}^{\prime},\theta^{\star}/\|\theta^{\star}\|\rangle\right|\leq 1/T^{3}\right)$
@@
$\mathbb{P}\left(\forall t,Y_{t}^{\prime}\langle X_{t}^{\prime},\theta^{\star}\rangle>0\right)$ $\geq\left(\left(1-\frac{\sqrt{d}}{\sqrt{\pi}T^{3}}\right)\cdot\frac{1}{1+e^{-T}}\right)^{T}$
@@
Whenever the data $\{(X_{t}^{\prime},Y_{t}^{\prime})\}_{t\in[T]}$ is separable, the hard margin SVM will, with high probability, output a classifier $\hat{\theta}$ where the disagreement region for the linear separators induced by $\hat{\theta}$ and $\theta^{\star}$ is only $O\left(\frac{d+\log T}{T}\right)$, according to the recent result of Bousquet et al. (2020). This is concretized by the following Lemma 32.
Lemma 32.
Let $T>1$ be an integer. For any ${\theta^{\star}}\in\mathbb{R}^{d}$, with $\|{\theta^{\star}}\|>T^{2}$, suppose the $\{X_{t}^{\prime}\}_{t=1}^{T}$ are drawn i.i.d. from the distribution $\mathcal{D}_{{\theta^{\star}}}$ conditioning on the event $\{Y^{\prime}=\operatorname{sign}(\langle X^{\prime},{\theta^{\star}}\rangle)\}$. Then, conditioning on the event $\{\forall t,Y_{t}^{\prime}=\operatorname{sign}(\langle X_{t}^{\prime},{\theta^{\star}}\rangle)\}$, with probability at least 1d/T1-d/T, the $\mathrm{SVM}$ solution $\hat{\theta}$ on $\{(X_{t}^{\prime},Y_{t}^{\prime})\}_{t\in[T]}$ achieves
$\hat{\theta}\in\operatorname*{arg\,min}_{\theta}\sum\limits_{t=1}^{T}\mathbbm{1}(Y_{t}^{\prime}\langle X_{t}^{\prime},\theta\rangle<0)$
@@
Proof of Lemma 32 If $d>T$, then the argument automatically holds. Thus, we assume hereafter that $d\leq T$. This proof is a revised version of Hsu and Mazumdar (2024, Lemma 14). For any $\theta$, the error of the linear separator defined by $\theta$ in $0/1$ loss is defined as
$\mathrm{err}_{{\theta^{\star}}}(\theta)$ $\vcentcolon=\mathbb{E}\left[\mathbbm{1}(Y\langle X,\theta\rangle<0)\right]$